# Risk-Calibrated Human-Robot Interaction via Set-Valued Intent Prediction

### v1

@@ Abstract
Tasks where robots must cooperate with humans, such as navigating around a cluttered home or sorting everyday items, are challenging because they exhibit a wide range of valid actions that lead to similar outcomes. Moreover, zero-shot cooperation between human-robot partners is an especially challenging problem because it requires the robot to infer and adapt on the fly to a latent human intent, which could vary significantly from human to human. Recently, deep learned motion prediction models have shown promising results in predicting human intent but are prone to being confidently incorrect. In this work, we present Risk-Calibrated Interactive Planning (RCIP), which is a framework for measuring and calibrating risk associated with uncertain action selection in human-robot cooperation, with the fundamental idea that the robot should ask for human clarification when unertainty in the human’s intent may adversely affect task performance. RCIP builds on the theory of set-valued risk calibration to provide a finite-sample statistical guarantee on the cumulative loss incurred by the robot while minimizing the cost of human clarification in complex multi-step settings. Our main insight is to frame the risk control problem as a sequence-level multi-hypothesis testing problem, allowing efficient calibration using a low-dimensional parameter that controls a pre-trained risk-aware policy. Experiments across a variety of simulated and real-world environments demonstrate RCIP’s ability to predict and adapt to a diverse set of dynamic human intents. [^1] Website with additional information, videos, and code: https://risk-calibrated-planning.github.io/
@@ I Introduction
Predicting and understanding human intent is a critical task for robotics, specifically for safe interaction with humans in cluttered, close-quarters environments. However, human intent prediction is challenging because no two humans may have the same preferences, and intents may differ depending on the specific environment. As an example, a robot is tasked with sorting items into three bins based on an example provided by the human (see Fig. 1). While the bins have a ground-truth sorting criterion known by the human (vegetables, children’s toys, and miscellaneous orange items), the robot must infer the human’s intent in order to sort new items. Given the provided context, the robot should be able to sort some unambiguous items (e.g. the crab) autonomously, while other items (e.g. the carrot) may be placed into multiple bins, resulting in situational ambiguity. If asked to operate fully autonomously, the robot must take a risk and guess the correct placement for the carrot. However, the robot may also ask for help if it is unsure, guaranteeing the correct action but potentially burdening the human. In this work, we study the tradeoff between risk and autonomy governing optimal action selection in the face of situational ambiguity.
@@
Our approach utilizes deep-learned human intent prediction models (e.g. [3, 4]) for understanding interactivity, and rigorously quantifies the uncertainty of these models in order to decide when to ask for help. As shown in Fig. 1 (middle), we produce a limited set of human intents based on the prediction model’s confidence scores. For each predicted intent, we plan a sequence of actions that satisfy an environment objective, such as placing the item in the correct bin. To accommodate different levels of robot autonomy, we assume that the predictor has a small number of highly flexible hyperparameters (such as the temperature), which allow the end-user to specify high-level behaviors (more or less confident predictions). We use a small calibration dataset of human-robot interactions to choose a set of valid hyperparameters that provide a level of risk and autonomy set in advance by the user. By leveraging recent advances in distribution-free risk control [5], we show that the robot’s behavior can simultaneously limit several notions of risk. We formalize this challenge via two objectives: (i) statistical risk calibration: the robot should seek sufficient help from the human when necessary to ensure a statistically guaranteed level risk specified by the user, and (ii) flexible autonomy: the robot should ask for a minimal amount of help as specified by the user by by narrowing down situational ambiguities through planning. We refer to these simultaneous objectives, with help from the human when necessary, as Risk-Calibrated Interactive Planning (RCIP).
@@
Statement of contributions. In this work, we introduce RCIP, a framework for measuring and calibrating risk in situations that involve interactions with humans with potentially ambiguous action choices. By reasoning about the human’s desired task outcome in the space of intents, we efficiently plan safe actions in the face of diverse, multi-modal human behavior, and ask for help when necessary. We make the following contributions: (1) We demonstrate how to use statistical risk control (SCR) to control the planning error rate across a set of model hyper-parameters, allowing flexible but provably safe levels of autonomy. (2) We prove theoretical guarantees for multi-dimensional risk control for both single-step and multi-step planning problems: with a set of user-specified risk budgets $(\alpha_{1},...,\alpha_{K})$ for different measures of risk (e.g., probability of failure and probability that the robot asks for help) and the robot performs the task correctly (with high probability) by asking for help if any of the $K$ risk budgets will be violated. (3) We evaluate RCIP in both simulation and hardware with a suite of human-robot interactive planning tasks with various styles of situational ambiguity (spatial, contextual, semantic). Experiments across multiple platforms and human uncertainty showcase the ability of RCIP to provide statistically guaranteed task success rates while providing more flexible autonomy levels than baseline approaches. RCIP reduces the amount of human help by $5-30\%$ versus baseline approaches.
@@ II-A Contingency Planning and Priviledged Learning
In this work, we provide the robot with additional information about the internal state of the “human” during the planning phase. We eliminate the need for a separate distillation procedure by instead using a set-valued prediction strategy, introduced in the following sections. We use contingency planning and privileged learning to find (or learn) optimal intent-conditioned policies, which can then be used to predict an optimal action via an upstream predictor. By allowing the robot to ask for help when it is uncertain, we statistically quantify risk associated with the robot acting optimally, even when it is uncertain.
@@ II-B Human Intent Prediction
Predicting intent of humans for downstream planning has been widely applied in autonomous driving [22, 23, 24], social navigation [25, 26, 4], and game theory [27]. Several works [23, 28] use a discrete latent variable to capture qualitative behaviors in human motion. To aid in human goal satisfaction, the authors of [29] show that human actions can be predicted directly in interactive settings, but the prediction model must be retrained whenever the task or human partner changes. Conversely, in our work we leverage recent advances in vision language models (VLMs) [30] for their ability to condition on internet-scale data to predict intuitive human motions in a variety of tasks. In this work, we use intent prediction to bound directly the risk associated with downstream planning. We use set-valued prediction to compute a set of possible intents, from which a planning module can compute a conditional plan.
@@ II-C Conformal Prediction and Empirical Risk Control
Conformal prediction [31, 32, 33] has recently gained popularity in a variety of machine learning and robotics applications due to its ability to rigorously quantify and calibrate uncertainty. A recent line of works [34, 35, 36] has extended the theory from prediction of labels (e.g. actions) to sequences (e.g. trajectories). Several works [37, 38] have studied adaptive conformal prediction, wherein a robot’s predictive conservativeness is dynamically adjusted within a policy rollout by assuming that there always exists a conservative fallback policy. Finally, some recent works [39, 40] have extended conformal prediction theory to handle more general notions of risk. Our work differs in three key ways: (i) we provide a separate calibration stage in which the robot can adjust its parameterization of prediction sets through a modest-size dataset of interactive scenarios, reducing the number of “unrecoverable” scenarios in which the robot exceeds its risk budget early on in a rollout, and (ii) we provide a way to synthesize from scratch risk-averse control policies, and (iii) we reason about human uncertainty in the space of intents, permitting a more natural way to capture diverse interactive behaviors than other representations (e.g. trajectories).
@@ III-A Dynamic Programming with Intent Uncertainty
Figure 2: RCIP formulates interactive planning as a multi-hypothesis risk control problem. Using a small set of calibration scenarios, RCIP computes step-wise prediction losses to form an aggregate emperical risk estimate. Using a risk limit, for each pair $(\lambda,\theta)$ of prediction thresholds and model hyperparameters, RCIP evaluates the hypothesis that the test set risk is above the limit. Thus, for all hypotheses that are rejected, the test set risk satisfies the threshold (with high probability).
@@
$x_{t+1}=f_{e}(x_{t},u_{t})\quad\forall t\in[T],$ (1)
where $x_{t}\in\mathcal{S}\subseteq\mathbb{R}^{n}$ is the joint state of the system and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the joint (robot-human) control input $(u_{t}^{R},u_{t}^{H})$. We use a superscript for individual agent indexing, and we use bar notation to denote aggregation over time, e.g. $\bar{x}_{t}=(x_{1},...,x_{t})$. Let $\pi=(\pi^{R},\pi^{H})$ denote the joint control policy governing system (1). We permit the human’s action to be drawn from a potentially multi-modal distribution $\pi^{H}$.
We present a methodology for learning a policy set $\Pi^{R}$ that selects a set of optimal actions contingent on the uncertainty in $\pi^{H}$.
@@
$z_{t+1}\sim q(\cdot|x_{t},z_{t}),$ (2)
where $z_{t}\in\mathcal{Z}=[N]$ characterizes the human’s intent at time $t$, and $N$ is the number of high-level human behaviors. We assume that conditioned on the human’s latent intent (which may be stochastic), each agent’s action is conditionally deterministic, i.e., $u_{t}^{i}=\pi^{i}(x_{t},z_{t})$, for $i\in\{R,H\}$.
@@
$J^{i}(x,z,\pi^{i})=\mathbb{E}^{\pi}\Bigg{[}\sum_{t=1}^{T}\;\;l^{i}(x_{t},u_{t})\Big{|}x_{1}=x,z_{1}=z\Bigg{]}.$ (3)
@@
Conditional Action Selection. For each intent $z$, the value function associated with the latent intent $z$ can be evaluated as a function of the state-intent pair $(x,z)$. That is,
@@
$Q^{i}(x,u^{i},z)=l^{i}(x,u^{i},u^{\neg i})+\mathbb{E}^{\pi,q}[V^{i}(x_{t+1},z_{t+1})],$ (7)
where $(x_{t+1},z_{t+1})$ are the next state-intent pair under the state dynamics (1) and intent dynamics (2), $\neg i$ is the other agent, and $u^{\neg i}=\pi^{\neg i}(x,z)$. Eqn. (7) states that since both agents’ policies are conditionally deterministic, the optimal cost-to-go is a deterministic function of the state-intent pair $(x,z)$. If the latent intent $z$ were known, then the expected cost-to-go is a deterministic function of the current observation $x$, intent $z$, and the joint control $u$. Without knowledge of $z$, the state of the system would not be fully observable, and the cost-to-go would vary due to uncertain action selection at the current state. The optimal action for both agents is given by the minimizer of Eqn. (7),
$u^{i*}(z)=\operatorname*{\arg\!\min}_{u}\;\;Q^{i}(x,u,z)\quad\forall i\in\{R,H\}.$ (8)
Intent Prediction. During deployment, the true human intent $z$ is not observed. However, we assume access to a model $g_{\theta}$ with hyperparameters $\theta$ that predicts human intent based on a sequence $\bar{x}_{t}$ of prior states. Specifically, let $g_{\theta}(\bar{x}_{t},z)$ produce a confidence score in $[0,1]$ that estimates the probability of each possible intent $z$ in $\mathcal{Z}$. In general, $g_{\theta}$ will output heuristic and uncalibrated confidence scores, and need not be trained on the same distribution $q$.
@@ III-B Risk-Calibrated Interactive Planning
Predicted Action Set. We aggregate confidence scores from $g_{\theta}$ into a set $\mathcal{S}_{\lambda,\theta}\subseteq\mathcal{Z}$ of predicted intents via the rule
@@
$\mathcal{T}_{\lambda,\theta}(\bar{x}_{t})=\{u\in\mathcal{U}:\exists z\in\mathcal{S}_{\lambda,\theta}\text{ s.t. }u=u^{*}(z)\text{ and }g_{\theta}^{*}(\bar{x}_{t},z)\geq\lambda\},$ (10)
where we define $g_{\theta}^{*}$ as the sum of all intent-based confidence scores that lead to the same action, i.e., $g_{\theta}^{*}(\bar{x}_{t},z):=\sum_{z^{\prime}\in\mathcal{Z}}g_{\theta}(\bar{x}_{t},z^{\prime})\mathbbm{1}\{u^{R*}(z)=u^{R*}(z^{\prime})\}$. To simplify notation, we define $g_{\theta}^{*}(z):=g_{\theta}^{*}(\bar{x}_{t},z)$.
@@
If $\lambda$ and $\theta$ are chosen such that $\mathcal{T}_{\lambda,\theta}(\bar{x}_{t})$ is empty, the task is failed.

### v2

@@ Abstract
Tasks where robots must anticipate human intent, such as navigating around a cluttered home or sorting everyday items, are challenging because they exhibit a wide range of valid actions that lead to similar outcomes. Moreover, zero-shot intent-prediction in human-robot cooperative scenarios is an especially challenging problem because it requires the robot to infer and adapt on the fly to a latent human intent, which could vary significantly from human to human. Recently, deep learned motion prediction models have shown promising results in predicting human intent but are prone to being confidently incorrect. In this work, we present Risk-Calibrated Interactive Planning (RCIP), which is a framework for measuring and calibrating risk associated with uncertain action selection in human-robot cooperation, with the fundamental idea that the robot should ask for human clarification when uncertainty in the human’s intent may adversely affect task performance. RCIP builds on the theory of set-valued risk calibration to provide a finite-sample statistical guarantee on the risk incurred by the robot while minimizing the cost of human clarification in complex multi-step settings. Our main insight is to frame the risk control problem as a sequence-level multi-hypothesis testing problem, allowing efficient calibration using a low-dimensional parameter that controls a pre-trained risk-aware policy. Experiments across a variety of simulated and real-world environments demonstrate RCIP’s ability to predict and adapt to a diverse set of dynamic human intents. [^1] Website with additional information, videos, and code: https://risk-calibrated-planning.github.io/
@@ I Introduction
Predicting and understanding human intent is a critical task for robotics, specifically for safe interaction with humans in cluttered, close-quarters environments. However, human intent prediction is challenging because no two humans may have the same preferences, and intents may differ depending on the specific environment. As an example, a robot is tasked with sorting items into three bins based on an example provided by the human (see Fig. 1). While the bins have a ground-truth sorting criterion known by the human (vegetables, children’s toys, and miscellaneous orange items), the robot must infer the human’s intent in order to sort new items. Given the provided context, the robot should be able to sort some unambiguous items (e.g., the crab) autonomously, while other items (e.g., the carrot) may be placed into multiple bins, resulting in situational ambiguity. If asked to operate fully autonomously, the robot must take a risk and guess the correct placement for the carrot. However, the robot may also ask for help if it is unsure, guaranteeing the correct action but potentially burdening the human. In this work, we study the tradeoff between risk and autonomy governing optimal action selection in the face of situational ambiguity.
@@
Our approach utilizes deep-learned human intent prediction models (e.g., [3, 4]) for understanding interactivity, and rigorously quantifies the uncertainty of these models in order to decide when to ask for help. As shown in Fig. 1 (middle), we produce a limited set of human intents based on the prediction model’s confidence scores. For each predicted intent, we plan a sequence of actions that satisfy an environment objective, such as placing the item in the correct bin. Depending on the robot’s confidence level and the human’s preferred level of autonomy, the robot can either take a risk or ask for help. To allow the human to specific different levels of robot autonomy (more or less confident predictions), we assume that the predictor has a small number of tunable model parameters (such as the temperature used in softmax scoring). We use a small calibration dataset of human-robot interactions to choose a set of valid parameters that provide a level of risk and autonomy set in advance by the user. By leveraging recent advances in distribution-free risk control [5], we show that the robot’s behavior can simultaneously limit several notions of risk. We formalize this challenge via two objectives: (i) statistical risk calibration (SRC): the robot should seek sufficient help from the human when necessary to ensure a statistically guaranteed level risk specified by the user, and (ii) flexible autonomy: the robot should ask for a minimal amount of help as specified by the user by narrowing down situational ambiguities through planning. We refer to these simultaneous objectives, with help from the human when necessary, as Risk-Calibrated Interactive Planning (RCIP).
@@
Statement of contributions. In this work, we introduce RCIP, a framework for measuring and calibrating risk in situations that involve interactions with humans with potentially ambiguous action choices. By reasoning about the human’s desired task outcome in the space of intents, we efficiently plan safe actions in the face of diverse, multi-modal human behavior, and ask for help when necessary. We make the following contributions: (1) We demonstrate how to use SRC to control the planning error rate across a set of model hyper-parameters, allowing flexible but provably safe levels of autonomy. (2) We prove theoretical guarantees for multi-dimensional risk control for both single-step and multi-step planning problems: with a set of $K$ user-specified risk budgets $(\alpha_{1},...,\alpha_{K})$ for different measures of risk (e.g., probability of failure and probability that the robot asks for help) and the robot performs the task correctly (with high probability) by asking for help if any of the $K$ risk budgets will be violated. (3) We evaluate RCIP in both simulation and hardware with a suite of human-robot interactive planning tasks with various styles of situational ambiguity (spatial, contextual, semantic). Experiments across multiple platforms and human uncertainty showcase the ability of RCIP to provide statistically guaranteed task success rates while providing more flexible autonomy levels than baseline approaches. RCIP reduces the amount of human help by $5-30\%$ versus baseline approaches.
@@ II-A Contingency Planning and Priviledged Learning
In this work, we provide the robot with additional information about the internal state of the human during the planning phase (for example, their preference for a particular bin as shown in Fig.  1). We eliminate the need for a separate distillation procedure by instead using a set-valued prediction strategy, introduced in the following sections. We use contingency planning (and privileged learning) to find (or learn) an intent-conditioned policy, which can then be used to predict an optimal action via an upstream predictor. By allowing the robot to ask for help when it is uncertain, we statistically quantify risk associated with the robot acting optimally, even when it is uncertain. In practice, we show that intent-condition planning permits a modular approach to planning in uncertain environments: the intent predictor allows the robot to discern high-level uncertainly in human preferences for task completion, while a low-level intent-conditioned planner allows the robot to ensures that an acceptable intent-conditioned policy exists.
@@ II-B Human Intent Prediction
Predicting intent of humans for downstream planning has been widely applied in autonomous driving [22, 23, 24], social navigation [25, 26, 3], and game theory [27]. Several works [23, 28] use a discrete latent variable to capture qualitative behaviors in human motion. To aid in human goal satisfaction, the authors of [29] show that human actions can be predicted directly in interactive settings, but the prediction model must be trained on a set of representative interactions. To showcase RCIP’s ability to enable zero-shot intent-conditioned planning we leverage recent advances in vision language models (VLMs) [30] to predict human intent by extracting task-relevant semantic features through vision, without fine-tuning on specific interactions.
In this work, we show that both task-specific (e.g., [3]) and zero-shot intent predictors (e.g., [4]) can be used to achieve a variety of robot autonomy levels when combined with RCIP. We enable efficient transfer to downstream planning by restricting the support of the predictor to a small set of human intents, allowing the model to focus on high-level uncertainties.
@@ II-C Conformal Prediction and Empirical Risk Control
Conformal prediction [31, 32, 33] has recently gained popularity in a variety of machine learning and robotics applications due to its ability to rigorously quantify and calibrate uncertainty. A recent line of works [34, 35, 36] has extended the theory from prediction of labels (e.g. actions) to sequences (e.g. trajectories). Several works [37, 38] have studied adaptive conformal prediction, wherein a robot’s predictive conservativeness is dynamically adjusted within a policy rollout by assuming that there always exists a conservative fallback policy. Finally, some recent works [39, 40] have extended conformal prediction theory to handle more general notions of risk.
Our work differs in three key ways: (i) we provide a separate calibration stage in which the robot can adjust its parameterization of prediction sets through a modest-size dataset of interactive scenarios, reducing the number of “unrecoverable” scenarios in which the robot exceeds its risk budget early on in a rollout, and (ii) we provide a way to synthesize from scratch risk-averse control policies, and (iii) we reason about human uncertainty in the space of intents, permitting a more natural way to capture diverse interactive behaviors than other representations (e.g. trajectories).
@@ III-A Dynamic Programming with Intent Uncertainty
Figure 2: RCIP formulates interactive planning as a multi-hypothesis risk control problem. Using a small set of calibration scenarios, RCIP computes step-wise prediction losses to form an aggregate emperical risk estimate. Using a risk limit, for each pair $(\lambda,\theta)$ of prediction thresholds and tunable model parameters, RCIP evaluates the hypothesis that the test set risk is above the limit. Thus, for all hypotheses that are rejected, the test set risk satisfies the threshold (with high probability).
@@
$x_{t+1}=f_{t}(x_{t},u_{t};\ e)\quad\forall t\in[T],$ (1)
where $x_{t}\in\mathcal{S}\subseteq\mathbb{R}^{n}$ is the joint state of the system and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the joint (robot-human) control input $(u_{t}^{R},\ u_{t}^{H})$. We use $[T]$ notation to denote the set of natural numbers up to $T$. We use a superscript for individual agent indexing, and we use bar notation to denote aggregation over time, e.g. $\bar{x}_{t}=(x_{1},\ \dotsc,\ x_{t})$. Let $\pi=(\pi^{R},\ \pi^{H})$ denote the joint policy governing system (1). We permit the human’s action to be drawn from a potentially multi-modal distribution $\pi^{H}$. We model the modes of $\pi^{H}$ as corresponding to different human intents, which may be time-varying, and their dynamics is described next.
@@
$z_{t+1}\sim q_{t}(\cdot|x_{t},z_{t};\ H)\quad\forall t\in[T],$ (2)
where $z_{t}\in\mathcal{Z}=[N]$ characterizes the human’s intent at time $t$, and $N$ is the number of high-level human behaviors. We assume that conditioned on the human’s latent intent (which may be stochastic), each agent’s action is drawn from a (potentially time-varying) probability distribution, i.e., $u_{t}^{i}=\pi^{i}_{t}(\cdot|x_{t},z_{t})$, for $i\in\{R,\ H\}$.
@@
$J^{i}(x,z,\pi^{i})=\mathbb{E}^{\pi,q}\Bigg{[}\sum_{t=1}^{T}\;\;l^{i}(x_{t},u_{t})\Big{|}x_{1}=x,z_{1}=z\Bigg{]}.$ (3)
@@
Intent-Conditioned Action Selection. For each intent $z$, the value function associated with the latent intent $z$ can be evaluated as a function of the state-intent pair $(x,z)$. That is,
@@
$Q^{i}(x,z,u^{i})=l^{i}(x,u)+\mathbb{E}^{(u,z^{\prime})\sim(\pi,q)}[V^{i}(f(x,u;\;e),z^{\prime})].$ (7)
Eqn. (7) states that since both agents’ policies are conditioned on the human’s intent (interpreted as a state constraint), the optimal action can be approximated with a variety of existing policy optimization or dynamic programming techniques, e.g., PPO [41], A-star graph search [42], or Rapidly Exploring Random Trees (RRT) [43]. In particular, the optimal action can be approximated for both agents as the minimizer of the $Q$-function, i.e.,
$u^{i*}(z)=\operatorname*{\arg\!\min}_{u}\;\;Q^{i}(x,z,u)\quad\forall i\in\{R,\ H\}.$ (8)
We emphasize that, in reality, the robot need not exactly actuate the optimal action as long as the intent constraint in (4) is satisfied.
Intent Prediction. During deployment, the true human intent $z$ is not observed. However, we assume access to a frozen model $g_{\theta}$ with a small number of tunable parameters $\theta$ that predicts human intent based on a sequence $\bar{x}_{t}$ of prior states. Specifically, $g_{\theta}(\bar{x}_{t},z)$ seeks to estimate the probability mass $q(z|x_{t-1},z_{t-1})$ of each intent $z\in\mathcal{Z}$ by producing a set of softmax confidence scores on the simplex $\Delta^{|\mathcal{Z}|-1}$. That is,
$g_{\theta}(\bar{x}_{t},z)=\frac{\exp(G_{\theta}(\bar{x}_{t},z))}{\sum_{z^{\prime}\in\mathcal{Z}}\exp(G_{\theta}(\bar{x}_{t},z^{\prime}))}\quad\forall z\in[\mathcal{Z}],$ (9)
where $G_{\theta}$ outputs the logit scores for each intent. In general, $g_{\theta}$ will output heuristic and uncalibrated confidence scores, and the approximation of the true intent dynamics $q$ may be poor, since $g_{\theta}$ need not be trained on the same human intent dynamics as given by (2). However, we show in the following analysis (c.f. Section IV) that by asking for help, the robot is able to overcome incorrect predictions through strategic selection of human interventions with SRC.
@@ III-B Risk-Calibrated Interactive Planning
RCIP seeks to distill a weighted set of actions using the scores from $g_{\theta}$ and the actions from the intent-conditioned planner. By using an action set during action prediction, RCIP ensures that a plan exists for each possible human intent. We now describe how the action set is obtained.
Predicted Action Set. We first aggregate confidence scores from $g_{\theta}$ into an intermediate set $\mathcal{S}_{\lambda,\theta}\subseteq\mathcal{Z}$ of predicted intents via the rule
@@
$\mathcal{T}_{\lambda,\theta}(\bar{x}_{t})=\{u\in\mathcal{U}:\exists z\in\mathcal{S}_{\lambda,\theta}\text{ s.t. }u=u^{R*}(z)\text{ and }g_{\theta}^{*}(\bar{x}_{t},z)\geq\lambda\},$ (11)
where we define $g_{\theta}^{*}$ as the sum of all intent-based confidence scores that lead to the same action, i.e., $g_{\theta}^{*}(\bar{x}_{t},z):=\sum_{z^{\prime}\in\mathcal{Z}}g_{\theta}(\bar{x}_{t},z^{\prime})\mathbbm{1}\{u^{R*}(z)=u^{R*}(z^{\prime})\}$.
Remark.
Eqns. (10) and (11) convey the fact that the mapping from intents to actions can be many-to-one. Taking Fig. 1 as an example, if the predicted intent set is augmented to include {Vegetables, Produce, Toys, Orange}, the predicted intent set will have $|\mathcal{S}_{\lambda,\theta}|=4$ intents for sorting, but the predicted action set will only have $|\mathcal{T}_{\lambda,\theta}|=3$ actions, since Vegetables and Produce are the same bin and the action is the same. Quantifying uncertainty in action space allows RCIP to ask for help only when the predicted actions are different.
@@
If $\lambda$ and $\theta$ are chosen such that $\mathcal{T}_{\lambda,\theta}(\bar{x}_{t})$ is empty, the robot ceases operation and the task is failed.